OpenStudy (anonymous):

A function is given. Determine the average rate of change of the function between the given values of the variable. g(x)=9/x+5; x=0, x=h

OpenStudy (tkhunny):

This, \(\dfrac{9}{x} + 5\), as you have written, or did you intend \(\dfrac{9}{x+5}\)? In any case: \(\dfrac{g(h) - g(0)}{h-0}\). Evaluate and ponder.

OpenStudy (anonymous):

I tried um, and got 45-9h+45/h(h+5)